OpenStudy (anonymous):

x^2+3=-4x how do u dolve this with the quadratic formula? Ive tryed 4 hours and keep geting the wrong answer...

OpenStudy (reemii):

put all terms on one side (the left one for example), and use the formula involving the discriminant \(\Delta\): the solutions are given by the formula \(-\frac{-b\pm\sqrt\Delta}{2a}\) (if the equation is \(ax^2+bx+c=0\))

OpenStudy (anonymous):

Whats the discriminant?

OpenStudy (reemii):

it's the quantity noted \(\Delta\) , the value is \(b^2-4ac\).

OpenStudy (reemii):

the correct fraction is \(\frac{-b\pm\sqrt{\Delta}}{2a}\) (i put a minus sign in the first message that shouldn't be there)

OpenStudy (anonymous):

Im gettting a triangle under that √, id that right?

OpenStudy (reemii):

yes. what is your \(\Delta\) ? if it's correct then you will most likely find the two solutions.

OpenStudy (reemii):

i don't know really what you did.. write \(x^2+3=-4x\) as \(x^2+4x+3=0\) and recognize the values of \(a,b,c\). Here \(a=1,b=4,c=3\).

OpenStudy (anonymous):

ok. I didnt pout abc in that order, one sec

OpenStudy (anonymous):

*put

OpenStudy (anonymous):

4^2-4(1)(3) right so far?

OpenStudy (reemii):

yes.

OpenStudy (anonymous):

3 and 1?

OpenStudy (reemii):

-1 and -3. i think. please check.

OpenStudy (anonymous):

ur right

OpenStudy (reemii):

yup, then you got the answer.

OpenStudy (precal):